OpenStudy (anonymous):

i need help with this can anybody help me If the amount of air in a spherical balloon is 288 cu. in., what is the radius of the balloon? answers 6in 7.3in 72in

OpenStudy (campbell_st):

ok... so the volume of a sphere is \[V = \frac{4}{3} \pi r^3\] you know V = 288 so \[288 = \frac{4}{3} \pi r^3\] now solve for r